Draw the shear and bending-moment diagrams for the beams and loadings.
24kN/m 64kN・m S250×2 12ml-12--12ml 40kN 40N 40kN 60 kN/m 35 km VITTTTIIIIIIIIIIII. 11||||いい 12.4m | 3.0m 4.6m_ 12.0m」
Answer #1

Calrelations CFD E Bмр : 24 Non XNO Y Ecquilibrion for - Px=0 CZEMA = 0 - Pq xH = 64 -48=0 - Ry= -28KN Pfly = c. Py+ Dy -48=0HOWN HOKN Eaulinian el fories lefy zo RBT Ri=705 dow 35kN/m Goynimi mmmmmm 13 DE 40X16+18979.3+ 40x4.6+39413 - 4ox 2 - R9x74